iuytyuioiuytyuiop:

You take a package to the local shipping company. They charge a fixed base cost of $5 per package plus an additional $0.33 per pound. If P represents the number of pounds of your package, and C is the total cost of shipping your package, which of these is the correct relationship between P and C?

iuytyuioiuytyuiop:

@Extrinix

iuytyuioiuytyuiop:

C=0.05P+33 P=0.33C+5 P=0.05C+33 C=0.33P+5

jimthompson5910:

Here's a similar question. I've just changed the numbers around. The company charges $7 per package plus an additional $0.65 per pound. If P is the number of pounds, and C is the total cost, then what is the equation connecting the two? The answer would be C = 0.65P + 7 This is because we are charged the base fee of $7 no matter what the weight is Then on top of this 7 dollars, we have an additional 0.65P dollars. We can think of it like this 1 pound = 0.65*1 dollars 2 pounds = 0.65*2 dollars 3 pounds = 0.65*3 dollars and so on In general P pounds = 0.65p dollars Again this is added onto the initial base fee of $7